OpenStudy (anonymous):

Using complete sentences, describe the characteristics and parts of sigma notation. Please also provide a unique example. (2 points)

OpenStudy (anonymous):

\[\sum _{n=0}^{N} a^n\] This Sum begin at index n=0 , ends at n=N

OpenStudy (anonymous):

I am little worried about "2 points"

OpenStudy (anonymous):

So if you look at the image above in the earlier post. n is your index. It starts at 0, but could start at any value. The final term is going to be N, which is an arbitrary number. Then you are taking a raised to n, incrementing one by one and summing the terms. Hence the name summation notation. Hope this helps. http://www.tutorsean.net

OpenStudy (across):

\[\leftarrow------------\sum\] It's the devil's trident.